OpenStudy (anonymous):

solve 0 = 3x^2 + 18x + 22 using quadratic formula

OpenStudy (dumbcow):

a = 3 b=18 c=22 \[x = \frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\]

OpenStudy (anonymous):

I get down to where I have to get the square root of 60 but 60 isn't a perfect square

OpenStudy (accessdenied):

Are there any perfect squares as factors of 60?

OpenStudy (anonymous):

4*15

OpenStudy (anonymous):

\[-18+/- \sqrt{60}\]

OpenStudy (accessdenied):

Yeah, so we can rewrite it as the sqrt(4*15), and then break it apart \[ \begin{split} \sqrt{4*15} &= \sqrt{4} \sqrt{15}\\ &= 2\sqrt{15}\\ \end{split} \]

OpenStudy (anonymous):

so to go along with the equation i posted above, the 2 would be added to the -18?

OpenStudy (accessdenied):

nope, since the radical is still there, we cannot add them together. However, since the denominator is 6 (divisible by 2) and the numerator's terms (18 and 2sqrt(15)) are divisible by 2, we can cancel a 2 in both the numerator and the denominator. \[ \frac{18 \pm 2\sqrt{15}}{6} = \frac{\cancel{2}(9 \pm \sqrt{15})}{\cancel2*3} \]

OpenStudy (accessdenied):

and since we're not asked to approximate or anything ,its sufficient to leave it as that. \[ x= \frac{9 \pm \sqrt{15}}{3} \]

OpenStudy (dumbcow):

what do you write to show the 2's being crossed out ?

OpenStudy (accessdenied):

oops, I should include the negative. my bad!

OpenStudy (anonymous):

what if i was asked ti find the roots? would that answer till be okay?

OpenStudy (anonymous):

*to

OpenStudy (accessdenied):

dumbcow, \cancel{#} i had to look it up to find out how not long ago. :P yes, since the roots are the x values that make the equation = 0.

OpenStudy (anonymous):

okay, thanks a lot

OpenStudy (accessdenied):

@dumbcow \[ \frac{18 \pm 2\sqrt{15}}{6} = \frac{\cancel{2}(9 \pm \sqrt{15})}{\cancel{2}*3} \] \frac{18 \pm 2\sqrt{15}}{6} = \frac{\cancel{2}(9 \pm \sqrt{15})}{\cancel{2}*3} \cancel{#}

OpenStudy (dumbcow):

thanks